OpenStudy (juan1857):

Find the possible points of inflection: (x^3) +243/x

OpenStudy (juan1857):

I have got the 2nd derivative but I don't think there is a point of inflection...

OpenStudy (juan1857):

yeah

OpenStudy (solomonzelman):

You don't need to graph it.

OpenStudy (solomonzelman):

You know what concavity is, right? (i.e. you know what concave up and concave down means, correct?)

OpenStudy (juan1857):

yes

OpenStudy (solomonzelman):

Very good.

OpenStudy (solomonzelman):

In general, when is the function \(f(x)\) concave down?

OpenStudy (juan1857):

when f''(x) is negative

OpenStudy (solomonzelman):

Yes, good.

OpenStudy (solomonzelman):

And when is this function \(f(x)\) concave up?

OpenStudy (juan1857):

when f''(x) is positive

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

So, (for any function \(f(x)\)), the points of inflection are the points where concavity of the function changes.

OpenStudy (juan1857):

okk

OpenStudy (solomonzelman):

i.e. \(\color{black}{\displaystyle x=c }\) is a point of inflection, IFF, \(\color{black}{\displaystyle f''(c)=0 }\) and if \(\color{black}{\displaystyle f''(x)}\) has different signs before \(x=c\) and after \(x=c\).

OpenStudy (juan1857):

yeah but I am having trouble finding when f''(x) =0

OpenStudy (solomonzelman):

Have you found \(f''(x)\)?

OpenStudy (juan1857):

yes

OpenStudy (solomonzelman):

Can you tell me what it is?

OpenStudy (juan1857):

sure

OpenStudy (juan1857):

\[6x +\frac{ 486 }{ x ^{3} }\]

OpenStudy (solomonzelman):

Yes, very good.

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle f''(x)=6x+\frac{486}{x^3} }\)

OpenStudy (solomonzelman):

Now, in order to find possible points of inflection, set \(\color{black}{\displaystyle f''(x)=0 }\).

OpenStudy (juan1857):

\[x ^{4}=-81\]

OpenStudy (juan1857):

thats as far as i got

OpenStudy (juan1857):

but wouldn't you need "i"

OpenStudy (solomonzelman):

(Why do I say possible, because if \(\color{black}{\displaystyle f''(c+\delta) }\) and \(\color{black}{\displaystyle f''(c-\delta) }\) are both negative or both positive, then, despite that \(\color{black}{\displaystyle f''(c)=0 }\), the point \(x=c\) is not an inflection, since the concavity of the function doesn't change.)

OpenStudy (solomonzelman):

(For sufficiently small \(\delta\). ... In other words, we are (or would be) testing the concavity near x=c, to see if the concavity indeed changes.)

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle f''(x)=0 }\) \(\color{black}{\displaystyle 6x+\frac{486}{x^3}=0 }\)

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \frac{6x^4}{x^3}+\frac{486}{x^3}=0 }\) \(\color{black}{\displaystyle \frac{6x^4+486}{x^3}=0 }\) \(\color{black}{\displaystyle x^4=81 }\) So, far you are correct in solving for x.

OpenStudy (solomonzelman):

oh excuse me, -81.

OpenStudy (solomonzelman):

Obviously, \(\color{black}{\displaystyle \nexists x\in \mathbb{R}: x^4<0}\)

OpenStudy (solomonzelman):

i.e. \(\color{black}{\displaystyle x^4=-81}\) doesn't have any real-number solutions.

OpenStudy (solomonzelman):

That means there isn't an \(x\) to satisfy \(f''(x)=0\). (we aren't considering exclusively imaginary numbers)

OpenStudy (juan1857):

um ok, so should I put on my paper that it isn't concave up or concave down at any interval?

OpenStudy (solomonzelman):

Well, if there isn't a point \(x=c\) such that \(f''(c)=0\), then you definitely don't have any points of inflection.

OpenStudy (juan1857):

can you help you justify why there is no points of inflection?

OpenStudy (juan1857):

help me*

OpenStudy (solomonzelman):

Yes, I can.

OpenStudy (solomonzelman):

There are no point of inflection because (in this case) \(f''(x)\ne 0\) (for any \(x\)).

OpenStudy (juan1857):

oh ok

OpenStudy (solomonzelman):

any questions /

OpenStudy (solomonzelman):

?

OpenStudy (juan1857):

so how would I find concavity?

OpenStudy (juan1857):

just graph the 2nd derivative?

OpenStudy (solomonzelman):

Concave down iff \(\color{black}{\displaystyle f''(x)<0 }\). Concave up iff \(\color{black}{\displaystyle f''(x)>0 }\).

OpenStudy (juan1857):

but how can i find concavity if i can't find when f"(x)=0?

OpenStudy (solomonzelman):

These two equations above, in fact do have a solution. \(\large (\)Maybe there is a confusion as to how can my function \(f(x)\) have a solution to \(f''(x)>0\), and have a solution to \(f''(x)<0\), BUT not have a solution to \(x=0\)? (Doesn't this directly contradict the definition of an inflection point, which is defined to be the point where concavity changes, and we have concavity of different signs, but we don't have any points of inflection?!! The answer to this is, that \(x=0\) would have been our point of inflection (and you would see this even better if you solved \(f''(x)>0\) and \(f''(x)<0\)), however, \(x=0\) is not part of the function's domain, because there is a vertical asymptote at \(x=0\).\(\large )\)

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle f''(x) =\frac{x^4+486}{x^3} }\) and \(\color{black}{\displaystyle (x^4+486)>0 }\) for all \(x\), so whether \(f''(x)\) is negative or positive is going to depend entirely on \(x^3\). Set, \(x^3>0\) and \(x^3<0\) to see where the function is concave up, and where the function is concave down.

OpenStudy (solomonzelman):

An equivalent to this set up, which is as well the solution, set \(x>0\) to see where the function is concave up, and set \(x<0\) to see where the function is concave down. (Why? Because \(x^3>0\) iff \(x>0\), and \(x^3<0\) iff \(x<0\).)

OpenStudy (solomonzelman):

.... and there are the intervals where the function is concave down and up ...

OpenStudy (solomonzelman):